True or false, and justify.

Any subinterval (a,b) ⊂ (0,1) contains a number from the set

D := {1/n | n ∈N}
Answer #1

false

let a =1/2 , b = 1

so interval (a,b)=(1/2,1) ⊂ (0,1)

but ∄ x ∈ D such that x ∉ (1/2,1)